Question 4 

Finally, there is a last consideration: 

 

a)     Selling the company: Considering that our plant produces annual cashflows of $240,000 increasing at a 0,5% annual, expected to last for many years, for how much should we value the company today if investors are willing to obtain a profitability of 3%?



1
Expert's answer
2021-05-11T12:59:43-0400

The value of a company today refers to the intrinsic value. Intrinsic value is the right price which is used by investors and financial analysts to ascertain the actual value.

Valuation of company considers the future growth rate and profitability percentage in order to assess the right value. However, the valuation gets varies with the actual cash flow and expected cash flow.


Annual cash flows = $240,000

Growth rate = 0.5%

Profitability percentage = 3%

Value of a company today"=\frac{Annual\space cash\space flows\times (1+growth\space rate)}{(profitability\space percentage-growth\space rate)}"


"=\frac{240,000\times(1+0.5\%)}{(3\%-0.5\%)}"


"=\frac{241,200}{0.025}"


"=\$9,648,000"